DickT16:

ALGEBRA 2: Aden buys 15 collectible coins per month. Caliana sells 5 coins from her collection of 100 each month. When will Aden have more coins than Caliana? Show your work.

Kotesh:

Let's say months = x when Aden will have same coins as Caliana. Then 15×x = 100 - 5×x So x=5 After five months , Aden coins count 75 And Caliana coins count 75. So after six months Aden will have more coins than Caliana.

DickT16:

is that how to do the correct way in algebra 2 though?

bryanferia:

The number of months it will take for Aden to have more coins than Caliana is greater than 5 months How to write and solve equation? Aden: Number of coins bought per month = 15 Caliana: Number of coins sold from her collection = 5 let x = number of months Aden = Caliana 15x > 100 - 5x collect like terms 15x + 5x > 100 20x > 100 x > 100/20 x > 5 months

bryanferia:

@Pizzat16 wrote:
is that how to do the correct way in algebra 2 though?
check mine out

DickT16:

YES! yours is correct, thank you man

bryanferia:

@Cookiet16 wrote:
YES! yours is correct, thank you man
my pleasure man good luck

DickT16:

Thx